OpenStudy (anonymous):

Simplify by reducing the index of the radical (Help? I don't understand how to do this!):

OpenStudy (anonymous):

\[\sqrt[12]{125x ^{3}}\]

hero (hero):

Hint: What if we re-write it this way: \(\sqrt[12]{5^3x^3}\) Then change the radical to a fraction to get this: \((5x)^{\frac{3}{12}}\)

hero (hero):

That fraction can also be simplified. After simplification, you can change the fraction back to a radical.

OpenStudy (anonymous):

So does that mean the radical would look like: \[\sqrt[4]{}\]

OpenStudy (anonymous):

Oops, \[\sqrt[4]{5x}\]

hero (hero):

Correct

OpenStudy (anonymous):

So that's it? Can you help me on one more example so I can fully understand?

hero (hero):

The only step you're probably confused on is this: \(a^xb^x\) = \((ab)^x\)

OpenStudy (anonymous):

what about this then? since it has two different variables and exponents: \[\sqrt[49]{y ^{21}z ^{14}}\]

OpenStudy (anonymous):

@Hero

hero (hero):

The variables don't really matter. If you notice, the example I showed you also has two variables.

hero (hero):

If anything, that one is even easier since 21/49 = 3/7 and 14/49 = 2/7

OpenStudy (anonymous):

So do you do each individually...like: \[y ^{21/49}\]

hero (hero):

Yup

OpenStudy (anonymous):

and \[z ^{14/49}\]

hero (hero):

But after you simplify the fraction, put it back in root form

OpenStudy (anonymous):

After you simplify those they go to \[\sqrt[7]{y ^{3}z ^{2}}\]

OpenStudy (anonymous):

?

hero (hero):

Looks like you have it all figured out.

OpenStudy (anonymous):

Thanks so much!!

hero (hero):

All the credit goes to you. I didn't have to show you much before you figured it out.

OpenStudy (anonymous):

Would you happen to know anything about factoring and simplifying algebraic expressions? I honestly can't understand it for the life of me.

hero (hero):

Yup, I know about those

OpenStudy (anonymous):

Do you mind helping me understand how to do an expression like this: \[(x+7)^{-1/5}+(x+7)^{-6/5}\]

hero (hero):

Well, the first thing you should note is that a negative expression has nothing to do with "negative". In this case, negative exponent means "inverse".

hero (hero):

So in other words, you need to write the inverse of \((x+7)^{1/5}\) for example

hero (hero):

The inverse of a number is simply 1 over that number. Therefore \(x+7)^{-1/5}\) = \(\large \frac{1}{(x+7)^{1/5}} \)

OpenStudy (anonymous):

Okay, so the other would be \[\frac{ 1 }{ (x+7)^{6/5} }\]

hero (hero):

Yes

OpenStudy (anonymous):

Okay, and then what do I do front there?

OpenStudy (anonymous):

*from

hero (hero):

Hang on a minute

hero (hero):

One thing you can do is this: Rewrite \(\large \frac{ 1 }{ (x+7)^{6/5} }\) as \(\large\frac{1}{\sqrt[5]{(x+7)^5(x+7)^1}}\)

hero (hero):

Let me know if you understand that or not.

OpenStudy (anonymous):

the 5 and the 1 equal 6 so it's like 6/5

hero (hero):

Okay. Further more, you can split the denominator to this: \(\large\frac{1}{\sqrt[5]{(x+7)^5} \sqrt[5]{(x+7)^1}}\)

hero (hero):

Do you know what \(\sqrt[5]{(x+7)^5}\) simplifies to?

OpenStudy (anonymous):

Would it just be \[\sqrt{(x+7)}\]

hero (hero):

actually the root cancels as well

hero (hero):

leaving just x + 7

OpenStudy (anonymous):

Ohhh, so what about the 1/(x+7)1/5?

hero (hero):

Basically here's what you're left with now: \(\large \frac{1}{(x+7)\sqrt[5]{(x+7)}} + \frac{1}{\sqrt[5]{(x+7)}}\) What do you think you should multiply the second fraction by to create the same denominator?

OpenStudy (anonymous):

1/(x+7)?

hero (hero):

You can only multiply fractions by equivalents of 1 in this case when you're trying to create the same denominator

hero (hero):

In other words, multiply by (x+7)/(x+7)

hero (hero):

I hope I didn't lose you.

OpenStudy (anonymous):

Sorry, I'm here.

hero (hero):

What's the matter?

OpenStudy (anonymous):

I went to get something. So, do I multiply the \[\sqrt[5]{(x+7)}\] by (x+7)?

hero (hero):

You multiply the entire second fraction by (x+7)/(x+7) You should be able to interpret what I mean by that.

hero (hero):

The demoninators of both fractions are required to be the same before combining them. However, the numerators are not required to be the same.

OpenStudy (anonymous):

Okay.

hero (hero):

Let me know what you end up with.

OpenStudy (anonymous):

Well now it's just a mix of numbers and I'm confused. I have: \[\frac{ 1 }{ (x+7)\sqrt[5]{(x+7)} } +\] the same thing (sorry lost connection

OpenStudy (anonymous):

(If I don' reply, I'm eating dinner. But I'll be back! So, you can continue)

hero (hero):

You should have ended up with \(\large \frac{ x+8 }{ (x+7)\sqrt[5]{(x+7)} } \) If you had multiplied the second fraction by (x+7)/(x+7), thats what you would have gotten.

hero (hero):

You can also simplify the denominator as well

hero (hero):

Here's the next step you would perform to do that: \(\large \frac{ x+8 }{ (x+7)^{5/5}(x+7)^{1/5} }\) Then afterwards, you would use the rules for multiplying exponents

hero (hero):

Eventually, you would end up with \(\large \frac{ x+8 }{ \sqrt[5]{(x+7)^6} }\) as you final simplified form

OpenStudy (anonymous):

Thank you so much for your help. I'll be sure to ask you if I have anymore hard questions!

hero (hero):

Yes, I'm still here

OpenStudy (anonymous):

(: Thank you.

OpenStudy (anonymous):

It was such a confusing question!

OpenStudy (anonymous):

I actually have another question, if you're up for helping.

hero (hero):

You should post it as a new question, then close it immediately that way no one interferes.

OpenStudy (anonymous):

Alright!